Mrs Kumar paid $114 for 4 kg of prawns and 3 kg of crabs. If she had bought 3 kg of prawns and 5 kg of crabs instead, she would have to pay $21 more. What was the cost of 1 kg of crab?
Found 2 solutions by zhagi17, macston:
Answer by zhagi17(16)
Answer by macston(5194)

.
P=price/kg of prawns; C=price/kg of crabs
.
4kg(P)+3kg(C)=$114 Multiply this by 3.
12kg(P)+9kg(C)=$342 Call this Equation 1.
3kg(P)+5kg(C)=$135 Multiply this by 4.
12kg(P)+20kg(C)=$540 Call this Equation 2.
.
Subtract Equation 1 from Equation 2:
. 12kg(P)+20kg(C)=$540
-(12kg(P)+ 9kg(C)=$342)
--------------------------------
11kg(C)=$198
.
C=$18/kg
.
ANSWER: One kilogram of crab cost $18.